three point charges each of value +q are placed on the three vertices of a square of side "a" meters .what is the magnitude of force on point charge of value -q coloumb placed on the centre of the square

Grade:12

2 Answers

Chetan Mandayam Nayakar
12 years ago

Consider four charges +q at the vertices of a square. From symmetry, the net force is zero. Now consider an additional charge 

-q at the fourth vertex. From the principle of superposition this charge distribution is equivalent to the three charge configuration. Force=(1/4πk)((q^2)/(a^2/2))=(1/4πk)(2q^2)/a^2 

sanatan sharma
12 years ago

F =( K x q2 ) √2 / a

please approve the answer